# Quantum-Brain: Quantum-Inspired Neural Network Approach to Vision-Brain Understanding

### v1

@@ 4 The Proposed Quantum-Brain Approach
This section presents the vision-brain understanding problem and the limitations of prior vision-brain understanding methods. Then, we introduce the quantum-inspired neural network based on the entanglement. Finally, we summarize the learning framework for vision-brain understanding. Fig. 2 illustrates an overview framework of our proposed quantum-inspired neural network approach to vision-brain understanding.
@@ 4.1 Input Modeling
Let $X_{\text{image}}\in\mathbb{R}^{H\times W\times 3}$ and $\mathbf{x}\in\mathbb{R}^{C}$ be an image and its corresponding fMRI voxels where $H$, $W$, and $C$ is the height and width of the image and the number of voxels in the fMRI. An fMRI voxel encoder is applied to extract the features $\mathbf{p}\in\mathbb{R}^{D}$ of x and learn to align the extracted features to the high-level features $\mathbf{t}\in\mathbb{R}^{D}$ of $X_{\text{image}}$ for the semantic understanding, where $D$ is the dimension of the semantic features. In this work, we use CLIP ViT-L/14 [49] to build the semantic feature space of the images since the CLIP image encoders are trained to maximize the similarity with text captions of the images.
@@ 4.2 Connectivity of Brain Voxels via Entanglement
Limitations of Self-Attention. In the Vision-Brain Understanding problem, self-attention in Transformers [73] is one of the conventional approaches to compute the correlations or connectivities between elements. In the scenario of fMRI voxels, the connectivities between brain voxels depend on the information processing function of each brain region. Each brain region is responsible for specific information and can connect with others. For example, a low-level information region that directly processes information from human perception can connect to specific high-level information regions to process more complex semantic information. A conventional self-attention model learns to find patterns of brain activities without realizing that the connectivities depend on specific positions. Hence, the self-attention model fails to learn the connectivities in the brain signals. Fig. 3 illustrates the comparison between the connectivity learnability of the self-attention method and the proposed quantum-inspired network.
@@ 4.3 Quantum-Inspired Voxel-Controlling Module
It is crucial to take account of the connectivities in the fMRI signal to explore the semantic information. The controlling operators in quantum computing can utilize the connectivities between fMRI voxels. In detail, for each fMRI voxel $x_{j}$, its value can be formulated into a quantum state $|\psi_{j}\rangle=\sqrt{1-x_{j}}|0\rangle+\sqrt{x_{j}}|1\rangle$. Then, the voxel $x_{j}$ can be controlled by the other voxel $x_{k}$ via a controlling operator $U_{\text{control}}^{(j,k)}$ defined as Eqn. (4).
@@
Moreover, to enhance the connectivity representation, we apply a Phase-Shifting module calibrating the voxel values. In detail, the quantum state $|\psi_{j}\rangle$ representing the fMRI voxel $x_{j}$ can be formulated as Eqn. (6).
@@ 4.4 Measurement-like Projection Module
While the quantum state $|\psi_{j,k}\rangle$ is represented in the 4-dimensional Hilbert space $\mathcal{H}^{4}$, the connectivity information can be reduced into 2-dimensional Hilbert space $\mathcal{H}^{2}$. This is because the value of $\sqrt{1-x_{j}}$ in $|\psi_{k}\rangle\otimes|\psi_{j}\rangle$ is redundant for the projection since the value of $\sqrt{x_{j}}$ is also in that state. In detail, the learnable basis state $|\phi_{j,k}\rangle$ can be defined as Eqn. (8).
$|\phi_{j,k}\rangle=\begin{pmatrix}0,1,0,W_{j,k}\end{pmatrix}^{\top}$ (8)
Then, derived from Eqn. (7), the connectivity function from $x_{k}$ to $x_{j}$ can be formulated as Eqn. (9).
$\begin{split}f(x_{j},x_{k})&=x_{j}+x_{j}x_{k}(W_{j,k}^{2}-1)\\ &\quad\quad+x_{j}a_{k}\cdot 2W_{j,k}\cos{(\theta_{0}^{(k)}-\theta_{1}^{(k)})}\end{split}$ (9)
@@
Figure 5: The fMRI-to-image comparison between Takagi & Nishimoto [66], MinD-Vis [8], MindEye [58], and our proposed network on the Subject 1 NSD dataset [1]. Compared to prior methods, the images reconstructed by our proposed method have more consistent structures and closer semantic meaning to the reference images.
@@ 4.5 High-Level Feature Contrastive Learning
For vision-brain understanding tasks, the features of the fMRI signals should be aligned with the high-level features of the corresponding images. Contrastive learning is an effective learning method that represents multiple modalities. It maximizes the cosine similarity for positive pairs while minimizing the similarity for negative pairs. Contrastive language-image pre-trained model (CLIP) [49] is a multimodal contrastive model aligning the features of the images and their captions in texts into the same feature space. Because of the same feature space mapping, the features of the images extracted by CLIP have high semantic information obtained from the corresponding text captions.
Following the CLIP learning procedure, we use the CLIP loss [49] to learn the high-level features of the fMRI signals. Given $N$ pairs of fMRI feature $\mathbf{p}_{i}$ and its corresponding image feature $\mathbf{t}_{i}$, the CLIP loss is defined as Eqn. (12).
$\begin{split}\mathcal{L}_{\text{CLIP}}&=\sum_{i=1}^{N}\left[\log\left(\frac{\exp\left({\mathbf{p}_{i}^{\top}\cdot\mathbf{t}_{i}}/{\tau}\right)}{\sum_{j=1}^{N}\exp\left({\mathbf{p}_{i}^{\top}\cdot\mathbf{t}_{j}}/{\tau}\right)}\right)\right]\\ &\quad\quad+\sum_{i=1}^{N}\left[\log\left(\frac{\exp\left({\mathbf{t}_{i}^{\top}\cdot\mathbf{p}_{i}}/{\tau}\right)}{\sum_{j=1}^{N}\exp\left({\mathbf{t}_{i}^{\top}\cdot\mathbf{p}_{j}}/{\tau}\right)}\right)\right]\end{split}$ (12)
@@ 5.1 Implementation Details
Our experiments use CLIP ViT-L/14 [49] as a semantic image feature extraction. The reference images are resized into the resolution of $224\times 224$. The quantum-inspired network includes 4 blocks of voxel connection module. The fMRI voxels are extracted into $257\times 768$ dimensional features, aligning to the $257\times 768$ dimensional hidden CLIP features similar to [43, 58, 74]. We use the CLIP loss with a temperature $\tau=4\times 10^{-3}$ for the vision-brain alignment and the pre-trained Versatile Diffusion model [76] for image reconstruction. For better fMRI-to-image reconstruction, we apply a diffusion prior [50] to align the output features with the input space of the pre-trained diffusion model. The learning rate is set to $3\times 10^{-4}$ with the Cosine learning rate scheduler [32]. The model is optimized using AdamW [31] with 240 epochs and a batch size of 32. The proposed Quantum-Brain model and experiments are implemented and performed in PyTorch [46] on a single NVIDIA A100 GPU.
Table 1: Quantitative comparison of fMRI-to-image reconstruction methods on the NSD benchmark. The bolded results indicate the best results and the underlined results indicate the secondary best results.
@@ 5.2 Datasets and Benchmarks
We use the Natural Scenes Dataset (NSD) [1] for evaluation. This public fMRI dataset consists of the brain responses of 8 participants observing natural images from the MS-COCO dataset [28]. We can study the semantic features of the brain activities via the natural images corresponding to the fMRI signals. Following [58, 43, 66], we train subject-specific models for 4 participants separately. For each participant, there are $25,962$ samples split into $24,980$ training samples and $982$ testing samples, similar to [58, 43, 66].
To evaluate the performance of the proposed quantum-inspired neural network, we perform two vision-brain understanding tasks including fMRI-to-image reconstruction and image-brain retrieval. To demonstrate the feature representation capacity of the proposed approach, we evaluate the two benchmarks without further fine-tuning. The evaluation protocols and experimental results will be described in the following sections.
@@
Fig. 5 illustrates the qualitative results of our proposed model in fMRI-to-image reconstruction compared to prior methods [66, 8, 58]. In addition to the high-quality images, our reconstructed images have a closer semantic meaning to the reference images compared to previous methods. We also provide quantitative results of the proposed approach as shown in Table 1. Following Ozcelik and VanRullen [43], we evaluate the methods with 8 different image quality metrics. PixCorr is the pixel-level correlation metric between reconstructed images and ground-truth images. SSIM is the structural similarity index metric. AlexNet(2) and AlexNet(5) are the 2-way comparisons of the second and fifth layers of AlexNet, respectively. Inception is the 2-way comparison of the last pooling layer of InceptionV3. CLIP is the 2-way comparison of the output layer of the CLIP-Vision model. EffNet-B and SwAV are distance metrics gathered from EfficientNet-B1 and SwAV-ResNet50 models. The first four are low-level metrics, while the last four express higher-level properties. As illustrated in Table 1, our proposed model extracts fMRI information and achieves better high-level similarities to the corresponding images compared to prior methods. Meanwhile, the proposed model obtains competitive low-level results. It shows that the features extracted by the proposed quantum-inspired neural network are well-aligned with the semantic feature space of the images.
Table 2: Image-Brain retrieval results on the NSD benchmark.
@@
Ours Quantum-Brain 95.1% 95.6%
Table 3: Effectiveness of our method on the NSD Image-Brain Retrieval.
@@
95.1% 95.6%
5.4 Image-Brain Retrieval
Image-brain retrieval benchmarks study the level of fine-grain vision information contained in the brain embeddings. Following Lin et al. [27], the image-brain retrieval benchmarks evaluate the models in two separate procedures. For the image retrieval task, each testing fMRI sample is extracted into an fMRI embedding. Then we compute the cosine similarity to the CLIP image embeddings of its corresponding image and 299 other random images. For each testing sample, a correct retrieval is determined when the cosine similarity of the fMRI embedding and its corresponding image embedding is the highest. We average the retrieval results across all testing samples and repeat the whole process 30 times, similar to [27]. For brain retrieval, the procedure is similar to the image retrieval except we randomly select 299 other fMRI samples.
As shown in Table 2, our model outperforms previous methods [27, 43, 58] by a large margin. On the image-brain retrieval benchmark, the proposed approach achieves higher top-1 accuracy of $95.1\%$ and $95.6\%$ on the image and brain retrieval evaluations, respectively. It shows that the proposed model has a better and more balanced performance on the image-brain retrieval benchmark.
5.5 Ablation Studies
Our ablation experiments study the effectiveness of our proposed modules on the NSD image-brain retrieval benchmark as shown in Table 3.
Effectiveness of Phase-Shifting Module. As shown in Table 3, the Phase-Shifting module calibrates and better represents the brain voxel values for later information extraction. In detail, the top-1 accuracies of image retrieval have been increased by $0.9\%$, $1.7\%$, and $1.1\%$ for the three settings. Moreover, the Phase-Shifting module helps the Voxel-Controlling module extract the brain information better when the increasing of the top-1 accuracies of the image and brain retrieval has been improved from $0.7\%$ to $1.5\%$ and from $0.7\%$ to $4.2\%$.
Effectiveness of Controlling Module. Table 3 illustrates the impact of the Voxel-Controlling module. As shown in our results, the Voxel-Controlling module has improved the top-1 accuracies of the image and brain retrieval from $90.7\%$ and $86.4\%$ to $91.4\%$ and $87.1\%$ without Phase-Shifting module, and from $91.6\%$ and $87.2\%$ to $93.1\%$ and $91.4\%$ with Phase-Shifting module.
Effectiveness of Measurement-like Projection Module. As reported in Table 3, the Measurement-like Projection Module helps to transform the information in the Hilbert space into the feature space. Hence, the extracted information can be well-aligned with the semantic information obtained from the image CLIP features. In particular, the top-1 accuracies of the image and brain retrieval have been increased by $2.6\%$ and $1.5\%$ without the Phase-Shifting module and by $2.0\%$ and $4.2\%$ with the Phase-Shifting module.
@@ 6 Conclusions
This paper has introduced Quantum-Brain, a new quantum-inspired neural network for vision-brain understanding. Motivated by the entanglement properties in quantum computing theory, we have proposed a novel Quantum-Inspired Voxel-Controlling module to compute the connectivities between fMRI voxels represented in the Hilbert space. Then, a new Phase-Shifting module has been introduced to make the Quantum-Inspired Voxel-Controlling module learn the connectivities more effectively. Finally, we have presented a novel Measurement-like Projection Module to transform the connectivity information from the Hilbert space to the feature space. Our experimental results have demonstrated the effectiveness and significance of our proposed quantum-inspired neural network.
@@
Limitations. This study used a specific network design and hyperparameters to support our hypothesis. Our experiments are limited to the standard scale of the benchmarks due to the computational limitation. However, our proposed approach can be generalized for larger-scale benchmarks based on our theoretical analysis.

### v2

@@ 4 The Proposed Quantum-Brain Approach
This section presents the vision-brain understanding problem and the limitations of prior vision-brain understanding methods. Then, we introduce the quantum-inspired neural network based on entanglement. Finally, we summarize the learning framework for vision-brain understanding. Fig. 2 illustrates an overview of the framework for our proposed quantum-inspired neural network approach to vision-brain understanding.
@@ 4.1 Input Modeling
Let $X_{\text{image}}\in\mathbb{R}^{H\times W\times 3}$ and $\mathbf{x}\in\mathbb{R}^{C}$ be an image and its corresponding fMRI voxels where $H$, $W$, and $C$ is the height and width of the image and the number of voxels in the fMRI. An fMRI voxel encoder is applied to extract the features $\mathbf{p}\in\mathbb{R}^{D}$ of x and learn to align the extracted features to the high-level features $\mathbf{t}\in\mathbb{R}^{D}$ of $X_{\text{image}}$ for the semantic understanding, where $D$ is the dimension of the semantic features. In this work, we utilize CLIP ViT-L/14 [44] to construct the semantic feature space of the images, as the CLIP image encoders are trained to maximize the similarity with text captions of the images.
@@ 4.2 Connectivity of Brain Voxels via Entanglement
Limitations of Self-Attention. In the Vision-Brain Understanding problem, self-attention in Transformers [89] is a conventional approach to computing correlations or connectivities between elements. In the scenario of fMRI voxels, the connectivities between brain voxels depend on the information processing function of each brain region. Each brain region is responsible for processing specific information and can connect with other regions. For example, a low-level information region that directly processes information from human perception can connect to specific high-level information regions to process more complex semantic information. A conventional self-attention model learns to find patterns of brain activities without realizing that the connectivities depend on specific positions. Hence, the self-attention model fails to learn the connectivity in the brain signals. Fig. 3 illustrates the comparison between the connectivity learnability of the self-attention method and the proposed quantum-inspired network.
@@ 4.3 Quantum-Inspired Voxel-Controlling Module
It is crucial to consider the connectivity in the fMRI signal to explore semantic information. The controlling operators in quantum computing can utilize the connectivity between fMRI voxels. In detail, for each fMRI voxel $x_{j}$, its value can be formulated into a quantum state $|\psi_{j}\rangle=\sqrt{1-x_{j}}|0\rangle+\sqrt{x_{j}}|1\rangle$. Then, the voxel $x_{j}$ can be controlled by the other voxel $x_{k}$ via a controlling operator $U_{\text{control}}^{(j,k)}$ defined as Eqn. (3).
@@
Moreover, to enhance the connectivity representation, we apply a Phase-Shifting module, calibrating the voxel values. In detail, the quantum state $|\psi_{j}\rangle$ representing the fMRI voxel $x_{j}$ can be formulated as Eqn. (5).
@@ 4.4 Measurement-like Projection Module
While the quantum state $|\psi_{j,k}\rangle$ is represented in the 4-dimensional Hilbert space $\mathcal{H}^{4}$, the connectivity information can be reduced to a 2-dimensional Hilbert space $\mathcal{H}^{2}$. This is because the value of $\sqrt{1-x_{j}}$ in $|\psi_{k}\rangle\otimes|\psi_{j}\rangle$ is redundant for the projection since the value of $\sqrt{x_{j}}$ is also in that state. In detail, the learnable basis state $|\phi_{j,k}\rangle$ can be defined as $|\phi_{j,k}\rangle=\begin{pmatrix}0,1,0,W_{j,k}\end{pmatrix}^{\top}$. Then, derived from Eqn. (6), the connectivity function from $x_{k}$ to $x_{j}$ can be formulated as Eqn. (7).
$\begin{split}f(x_{j},x_{k})=x_{j}&+x_{j}x_{k}(W_{j,k}^{2}-1)\\ &+x_{j}a_{k}\cdot 2W_{j,k}\cos{(\theta_{0}^{(k)}-\theta_{1}^{(k)})}\end{split}$ (7)
@@
Figure 5: The diagram of the fMRI feature extraction.
Figure 6: The fMRI-to-image comparison between Takagi & Nishimoto [12], MinD-Vis [20], MindEye [13], and our proposed network on the Subject 1 NSD dataset [34]. Compared to prior methods, the images reconstructed by our proposed method exhibit more consistent structures and a closer semantic resemblance to the reference images.
5 fMRI Feature Extraction
As shown in Figure 5, the fMRI voxels are flattened into a voxel vector $\mathbf{x}^{(1)}\in\mathbb{R}^{C}$ where $C$ is the number of voxels. The number of voxels varies between different subjects from approximately 13,000 to 16,000 voxels. The voxel vector $\mathbf{x}^{(1)}$ is computed via connectivity functions $f$ defined as Eqn. (9). The computed voxel vector $\mathbf{x}^{(2)}$ is then extracted via a linear layer followed by two residual blocks with a hidden dimension of $4096$. The feature result of the final block is projected into a semantic feature $\mathbf{p}\in\mathbb{R}^{D}$. The feature dimension is denoted as $D=257\times 768$ to match the input feature of the Versatile Diffusion model.
@@ 5.1 High-Level Feature Contrastive Learning
For vision-brain understanding tasks, the features of the fMRI signals should be aligned with the high-level features of the corresponding images. Contrastive learning is an effective learning method that represents multiple modalities. It maximizes the cosine similarity for positive pairs while minimizing the similarity for negative pairs. The Contrastive Language-Image Pre-trained Model (CLIP) [44] is a multimodal contrastive model that aligns the features of images and their captions in texts into the same feature space. Due to the same feature space mapping, the features extracted from images by CLIP contain high semantic information corresponding to the text captions.
Following the CLIP learning procedure, we use the CLIP loss [44] to learn the high-level features of the fMRI signals. Given NN pairs of fMRI feature $\mathbf{p}_{i}$ and its corresponding image feature $\mathbf{t}_{i}$, the CLIP loss is defined as Eqn. (10).
$\begin{split}\mathcal{L}_{\text{CLIP}}&=\sum_{i=1}^{N}\left[\log\left(\frac{\exp\left({\mathbf{p}_{i}^{\top}\cdot\mathbf{t}_{i}}/{\tau}\right)}{\sum_{j=1}^{N}\exp\left({\mathbf{p}_{i}^{\top}\cdot\mathbf{t}_{j}}/{\tau}\right)}\right)\right]\\ &\quad+\sum_{i=1}^{N}\left[\log\left(\frac{\exp\left({\mathbf{t}_{i}^{\top}\cdot\mathbf{p}_{i}}/{\tau}\right)}{\sum_{j=1}^{N}\exp\left({\mathbf{t}_{i}^{\top}\cdot\mathbf{p}_{j}}/{\tau}\right)}\right)\right]\end{split}$ (10)
@@ 6.1 Implementation Details
Our experiments use CLIP ViT-L/14 [44] as a semantic image feature extraction. The reference images are resized to a resolution of $224\times 224$. The quantum-inspired network includes a voxel connection module as shown in Eqn. (9) and a multi-layer perceptron. The fMRI voxels are extracted into $257\times 768$ dimensional features, aligning to the $257\times 768$ dimensional hidden CLIP features similar to [18, 13, 14]. We use the CLIP loss with a temperature $\tau=4\times 10^{-3}$ for the vision-brain alignment and the pre-trained Versatile Diffusion model [10] for image reconstruction. For better fMRI-to-image reconstruction, we apply a diffusion prior [90] to align the output features with the input space of the pre-trained diffusion model. The learning rate is set to $3\times 10^{-4}$ with the Cosine learning rate scheduler [91]. The model is optimized using AdamW [92] with 240 epochs and a batch size of 32. The proposed model and experiments are implemented and performed in PyTorch [93] on a single NVIDIA A100 GPU.
TABLE I: Quantitative comparison of fMRI-to-image reconstruction on the NSD benchmark. The bolded results indicate the best results, and the underlined results indicate the secondary best results.
@@ 6.2 Datasets and Benchmarks
We use the Natural Scenes Dataset (NSD) [34] for evaluation. This public fMRI dataset comprises the brain responses of 8 participants viewing natural images from the MS-COCO dataset [95]. We can study the semantic features of brain activities through natural images corresponding to the fMRI signals. Following [13, 18, 12], we train subject-specific models for each of the four participants separately. For each participant, there are $25,962$ samples split into $24,980$ training samples and $982$ testing samples, similar to [13, 18, 12].
To evaluate the performance of the proposed quantum-inspired neural network, we perform two vision-brain understanding tasks, including fMRI-to-image reconstruction and image-brain retrieval. The evaluation protocols and experimental results will be described in the following sections.
@@
Fig. 6 illustrates the qualitative results of our proposed model in fMRI-to-image reconstruction compared to prior methods [12, 20, 13]. In addition to the high-quality images, our reconstructed images have a closer semantic meaning to the reference images compared to previous methods. We also provide quantitative results of the proposed approach as shown in Table I. Following Ozcelik and VanRullen [18], we evaluate the methods with eight different image quality metrics. PixCorr is a pixel-level correlation metric that measures the similarity between reconstructed images and ground-truth images. SSIM is the structural similarity index metric. AlexNet(2) and AlexNet(5) are the 2-way comparisons of the second and fifth layers of AlexNet, respectively. Inception is a two-way comparison of the last pooling layer in InceptionV3. CLIP is the 2-way comparison of the output layer of the CLIP-Vision model. EffNet-B and SwAV are distance metrics gathered from EfficientNet-B1 and SwAV-ResNet50 models. The first four are low-level metrics, while the last four express higher-level properties. As illustrated in Table I, our proposed model extracts fMRI information and achieves better high-level similarities to the corresponding images compared to prior methods. Meanwhile, the proposed model obtains competitive low-level results. It shows that the features extracted by the proposed quantum-inspired neural network are well-aligned with the semantic feature space of the images.
6.4 Image-Brain Retrieval
Image-brain retrieval benchmarks assess the level of fine-grained visual information contained in brain embeddings. Following Lin et al. [17], the image-brain retrieval benchmarks evaluate the models in two separate procedures. For the image retrieval task, each testing fMRI sample is extracted into an fMRI embedding. Then we compute the cosine similarity to the CLIP image embeddings of its corresponding image and 299 other random images. For each testing sample, a correct retrieval is determined when the cosine similarity of the fMRI embedding and its corresponding image embedding is the highest. We average the retrieval results across all testing samples and repeat the whole process 30 times. For brain retrieval, the procedure is similar to the image retrieval, except we randomly select 299 other fMRI samples.
TABLE II: Image-Brain retrieval results on the NSD benchmark.
@@ 6.4 Image-Brain Retrieval
Ours Quantum-Brain 95.5% 95.3%
As shown in Table II, our model outperforms previous methods [17, 18, 13] by a large margin. On the image-brain retrieval benchmark, the proposed approach achieves higher top-1 accuracy of $95.5\%$ and $95.3\%$ on the image and brain retrieval evaluations, respectively. The results show that the proposed model exhibits better and more balanced performance on the image-brain retrieval benchmark.
6.5 Ablation Studies
TABLE III: Effectiveness of our method on the NSD Image-Brain Retrieval.
Phase-
Shifting
Voxel-
Controlling
Measurement-like
Projection
@@ 6.5 Ablation Studies
95.5% 95.3%
Our ablation experiments study the effectiveness of our proposed modules on the NSD image-brain retrieval benchmark, as shown in Table III.
Effectiveness of Phase-Shifting Module. As shown in Table III, the Phase-Shifting module calibrates and better represents the brain voxel values for later information extraction. In detail, the top-1 accuracies of image retrieval have been increased by $0.9\%$, $1.7\%$, and $1.5\%$ for the three settings. Moreover, the Phase-Shifting module enables the Voxel-Controlling module to extract brain information more effectively, resulting in an increase in the top-1 accuracy of image and brain retrieval from $0.7\%$ to $1.5\%$ and from $0.7\%$ to $4.2\%$.
Effectiveness of Controlling Module. Table III illustrates the impact of the Voxel-Controlling module. As shown in our results, the Voxel-Controlling module has improved the top-1 accuracies of image and brain retrieval from $90.7\%$ and $86.4\%$ to $91.4\%$ and $87.1\%$, respectively, without the Phase-Shifting module, and from $91.6\%$ and $87.2\%$ to $93.1\%$ and $91.4\%$ with the Phase-Shifting module.
Effectiveness of Measurement-like Projection Module. As reported in Table III, the Measurement-like Projection Module helps to transform the information in the Hilbert space into the feature space. Hence, the extracted information can be well-aligned with the semantic information obtained from the image CLIP features. In particular, the top-1 accuracies of image and brain retrieval have increased by $2.6\%$ and $1.5\%$ without the Phase-Shifting module, and by $2.4\%$ and $3.9\%$ with the Phase-Shifting module.
@@ 7 Conclusions
This work has introduced Quantum-Brain, a novel quantum-inspired neural network for vision-brain understanding. Motivated by the entanglement properties in quantum computing theory, we have proposed a novel Quantum-Inspired Voxel-Controlling module to compute the connectivities between fMRI voxels represented in the Hilbert space. Then, a new Phase-Shifting module has been introduced to enable the Quantum-Inspired Voxel-Controlling module to learn connectivities more effectively. Finally, we have presented a novel Measurement-like Projection Module to transform the connectivity information from the Hilbert space to the feature space. Our experimental results have demonstrated the effectiveness and significance of our proposed quantum-inspired neural network.
@@
Limitations: This study used a specific network design and hyperparameters to support our hypothesis. Our experiments are limited to the standard scale of the benchmarks due to computational limitations. However, our proposed approach can be generalized for larger-scale benchmarks based on our theoretical analysis.
Broader Impacts: Our contributions open a novel paradigm for modeling complex brain connectivities, leveraging principles such as entanglement and superposition to enhance brain signal interpretation, and revolutionize the development of brain-computer interfaces by bridging theoretical quantum models with practical neuroscience applications.
Appendix A Expansion Details of Equation (7)
From Eqn. (4), the quantum state $|\psi_{j,k}\rangle$ can be represented as Eqn. (11).
$|\psi_{j,k}\rangle=\begin{pmatrix}\sqrt{(1-x_{k})(1-x_{j})}e^{i(\theta^{(k)}_{0}+\theta^{(j)}_{0})}\\ \sqrt{(1-x_{k})x_{j}}e^{i(\theta^{(k)}_{0}+\theta^{(j)}_{1})}\\ \sqrt{x_{k}(1-x_{j})}e^{i(\theta^{(k)}_{1}+\theta^{(j)}_{0})}\\ \sqrt{x_{k}x_{j}}e^{i(\theta^{(k)}_{1}+\theta^{(j)}_{1})}\end{pmatrix}$ (11)
Then the inner product between the basis state $|\phi_{j,k}\rangle$ and $|\psi_{j,k}\rangle$ is computed as Eqn. (12).
$\begin{split}\langle\phi_{j,k}|\psi_{j,k}\rangle&=\sqrt{(1-x_{k})x_{j}}e^{i(\theta^{(k)}_{0}+\theta^{(j)}_{1})}\\ &\quad\quad+W_{j,k}\sqrt{x_{k}x_{j}}e^{i(\theta^{(k)}_{1}+\theta^{(j)}_{1})}\end{split}$ (12)
Respectively, the inner product $\langle\psi_{j,k}|\phi_{j,k}\rangle$ is computed as follows,
$\begin{split}\langle\psi_{j,k}|\phi_{j,k}\rangle&=\langle\phi_{j,k}|\psi_{j,k}\rangle^{*}\\ &=\sqrt{(1-x_{k})x_{j}}e^{-i(\theta^{(k)}_{0}+\theta^{(j)}_{1})}\\ &\quad\quad+W_{j,k}\sqrt{x_{k}x_{j}}e^{-i(\theta^{(k)}_{1}+\theta^{(j)}_{1})}\end{split}$ (13)
where $\alpha^{*}$ indicates the complex conjugate of $\alpha$. Next, the connectivity function from $x_{k}$ to $x_{j}$ can be formulated as follows,
$\begin{split}f(x_{j},x_{k})&=f_{\text{MPM}}(|\psi_{j,k}\rangle,|\phi_{j,k}\rangle)\\ &=\langle\psi_{j,k}|\phi_{j,k}\rangle\langle\phi_{j,k}|\psi_{j,k}\rangle\\ &=(1-x_{k})x_{j}+W_{j,k}^{2}x_{k}x_{j}\\ &\quad\quad+W_{j,k}x_{j}\sqrt{(1-x_{k})x_{k}}\\ &\quad\quad\quad\cdot\left(e^{i(\theta^{(k)}_{0}-\theta^{(k)}_{1})}+e^{-i(\theta^{(k)}_{0}-\theta^{(k)}_{1})}\right)\\ &=x_{j}+x_{j}x_{k}(W_{j,k}^{2}-1)\\ &\quad+2W_{j,k}x_{j}\sqrt{(1-x_{k})x_{k}}\cos{(\theta^{(k)}_{0}-\theta^{(k)}_{1})}\end{split}$ (14)
since $e^{i\beta}+e^{-i\beta}=2\cos{(\beta)}$ from Euler’s formula. Then, we can define $a_{k}=\sqrt{(1-x_{k})x_{k}}$ as shown in Eqn. (7).
Appendix B fMRI Responses Computing
Quantum-Inspired Responses Computing. To compute the quantum-inspired responses from the V1 region to other regions as illustrated in Fig. 1 and Fig. 3, we drop out the values of all voxels except V1 as shown in Eqn. (15).
$\mathbf{x}^{\prime}=\mathbf{b}_{\text{mask}}\odot\mathbf{x}$ (15)
where $\mathbf{b}_{\text{mask}}$ is a masked vector of the V1 region, i.e., $\mathbf{b}_{\text{mask}}^{(i)}=1$ if the voxel $x_{i}$ is in the V1 region and otherwise $\mathbf{b}_{\text{mask}}^{(i)}=0$. Then, we utilize the Eqn. (9) to compute the responses $f(\mathbf{x}^{\prime})$ as shown in Eqn. (16).
$\begin{split}f(\mathbf{x}^{\prime})&=\mathbf{x}+\mathbf{x}\odot\left(\mathbf{W}^{\prime}\mathbf{x}^{\prime}\right)\\ &\quad\quad+\mathbf{x}\odot\left(\mathbf{W}^{\prime\prime}(\mathbf{a}^{\prime}\odot\cos{(\boldsymbol{\theta}_{0}-\boldsymbol{\theta}_{1})})\right)\end{split}$ (16)
where $\mathbf{a}^{\prime}=\sqrt{(1-\mathbf{x}^{\prime})\odot\mathbf{x}^{\prime}}$.
Attention Map for Responses Computing. To compute the attention map as shown in Fig. 1 and Fig. 3, we compute the attention scores obtained from the last layer of the Transformer model. Since we want to illustrate the connectivities from region V1 to other regions, we average the attention of the voxels in region V1 only to obtain the attention result.
Figure 7: The NSD fMRI-to-image reconstruction examples of the four subjects.
TABLE IV: Quantitative results of fMRI-to-image reconstruction on the NSD benchmark with different subjects.
  Subject Low-level High-level Retrieval
[…]
Subj01 .372 .361 96.9% 98.8% 96.7% 96.7% .591 .331 98.6% 97.4%
Subj02 .304 .356 97.1% 98.7% 95.2% 95.8% .625 .337 97.7% 97.6%
Subj05 .270 .331 93.2% 97.1% 96.6% 97.6% .584 .333 93.9% 93.7%
Subj07 .286 .323 94.1% 96.9% 93.7% 96.5% .643 .371 91.6% 92.5%
Appendix C Regions of Interest
In the Natural Scene Dataset, we utilize three regions of interest (ROIs), i.e., faces, places, and bodies, to visualize the performance of our model. These ROIs are obtained via functional localizer experiments (fLoc) 111 In the NSD dataset, this information can be obtained in nsddata/ppdata/subjAA/*/roi/, where AA is the name of the subject. . Note that these regions are not utilized for the training process.
Appendix D Subject-Specific Results
Fig. 7 shows the fMRI-to-image reconstruction examples of the four subjects 1, subject 2, subject 5, and subject 7. Additionally, Table IV reports the fMRI-to-image reconstruction and image-brain retrieval results of individual subjects.